OpenStudy (anonymous):

What is the average acceleration of a jet ski that goes from rest to 10 m/s west in 4 s? Take the direction west to be positive, and east to be negative. A. 2.5 m/s2 B. -2.5 m/s2 C. 0.4 m/s2 D. -0.4 m/s2

OpenStudy (jhannybean):

\[a_{avg} =\frac{\Delta v}{\Delta t}\]

OpenStudy (jhannybean):

that means if it starts at rest, \(v_i = 0 \ m/s\) while \(v_f= 10 \ m/s\) and the change in time is when \(t_i = 0 \ sec\) and \(t_f= 4\ sec\)

OpenStudy (jhannybean):

\[a_{avg} = \frac{\Delta v}{\Delta t} = \frac{v_f - v_i}{t_f -t_i} = \frac{10 \ m/s - 0 \ m/s}{4 \ sec - 0 \ sec}\]